(e none of the above answers (6 pts 7. In an R-L-C ac series circuit the voltage across the capacitor is 64 V and the voltage amplitude across the inductor is 96 V. Does the source voltage lag or lead the current in the circuit?can you explain how to answer this question
Answer #1

let I is the the current through the ckt.

Ic = IL = I (because they are in series)


we know,
VC = I*XC

VL = I*XL

so,

XL/XC = VL/VC

XL = (96/64)*XC

XL = 1.5*XC

so, XL > XC

we know phase angle, theta = tan^-1( (XL-XC)/R)

here theta is positive.

It means voltage leads the current in the ckt.